# Intelligent-Based Neural Networks and Optimal Control of Fractional Order Ebola Virus Dynamics

### v1

@@ 4.10 Global Stability of the Endemic Equilibrium
The fractional-order Ebola system (2) experiences a forward bifurcation at $\mathcal{R}_{0}=1$, without showing a backward bifurcation and without exhibiting multiple endemic equilibria.
Proof.  Global stability properties—disease-free equilibrium whenever $\mathcal{R}_{0}\leq 1$ and unique endemic equilibrium when $\mathcal{R}_{0}>1$—contradict the possibility of backward bifurcation. The epidemiological dynamics displays classical threshold behavior: disease elimination befalls when $\mathcal{R}_{0}\leq 1$ and $\mathcal{R}_{0}>1$ indicates endemic tenacity. $\square$
5 Sensitivity Analysis
5.1 Global Sensitivity Analysis and Parameter Influence
From the data, Latin Hypercube Sampling (LHS) estimates $\mathcal{R}_{0}$ and other critical epidemiological characteristics and examines the correlation of the parameters using PRCC. The PRCC examined global sensitivity for 10,000 samples within the reasonable biological boundaries. The outcomes are summarized in Table 3.
Table 3: Global sensitivity analysis: Parameter ranges and PRCC values for $\mathcal{R}_{0}$ and peak infection dynamics
Parameter Lower Bound Upper Bound PRCC ($\mathcal{R}_{0}$) PRCC (Peak $I_{s}$) p-value
$\beta$ 0.15 0.35 0.92 0.88 $p<0.001$
$\sigma$ 1/12 1/8 0.85 0.81 $p<0.001$
$\eta_{d}$ 0.4 0.9 0.68 0.65 $p<0.001$
pp 0.6 0.85 0.52 0.49 $p<0.001$
$\eta_{a}$ 0.3 0.7 0.41 0.38 $p<0.001$
$\varepsilon$ 0.7 0.95 -0.61 -0.58 $p<0.001$
$\gamma_{s}$ 1/16 1/12 -0.45 -0.42 $p<0.001$
$v$ 0.01 0.08 -0.35 -0.32 $p<0.001$
$\mu_{d}$ 0.1 0.3 -0.28 -0.25 $p<0.001$
$\alpha$ 0.7 0.95 -0.25 -0.23 $p<0.001$
$h_{s}$ 0.2 0.6 -0.21 -0.19 $p<0.001$
$\delta_{s}$ 0.05 0.12 0.15 0.18 $p=0.002$
5.2 Analytical Sensitivity Framework
The normalized sensitivity index quantifying $\mathcal{R}_{0}$ dependence on parameter $p$ is defined as:
$\Upsilon_{p}^{\mathcal{R}_{0}}=\frac{\partial\mathcal{R}_{0}}{\partial p}\times\frac{p}{\mathcal{R}_{0}}$ (18)
Theorem 5.1 (Analytical Sensitivity Indices)
The sensitivity indices for dominant parameters are given by:
Proof.  Analytical differentiation of the reproduction number yields.
$\text{Sensitivity indices for }\mathcal{R}_{0}:$ (19)
$\Upsilon_{\beta}^{\mathcal{R}_{0}}=+1$
$\Upsilon_{\sigma}^{\mathcal{R}_{0}}=1+\frac{\mathcal{R}_{0}^{(1)}}{\mathcal{R}_{0}}$
$\Upsilon_{\eta_{a}}^{\mathcal{R}_{0}}=\frac{\mathcal{R}_{0}^{(I_{a})}}{\mathcal{R}_{0}}$
$\Upsilon_{\eta_{d}}^{\mathcal{R}_{0}}=\frac{\mathcal{R}_{0}^{(D)}}{\mathcal{R}_{0}}$
$\Upsilon_{\varepsilon}^{\mathcal{R}_{0}}=-\frac{V^{*}}{S^{*}+(1-\varepsilon)V^{*}}$
$\Upsilon_{v}^{\mathcal{R}_{0}}=\frac{\beta\sigma}{\mathcal{R}_{0}(\Lambda/\mu)q_{2}}\cdot\frac{\partial}{\partial v}[S^{*}+(1-\varepsilon)V^{*}]\cdot K,$
The following equations illustrate the different components that contribute to the basic reproduction number:
$\mathcal{R}_{0}^{(I_{a})}$ $=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\cdot\frac{\eta_{a}(1-p)}{q_{4}}$ (20)
$\mathcal{R}_{0}^{(D)}$ $=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\cdot\frac{p\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}$ (21)
$\frac{\partial}{\partial v}[S^{*}+(1-\varepsilon)V^{*}]$ $=-\frac{\Lambda\varepsilon(\omega+\mu)}{\mu(q_{1}+v)^{2}}$ (22)
with the composite transmission factor given by
$K=\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}.$ (23)
From equations (20) and (21) one may discern the individual contributions of asymptomatic infections and post-mortem transmission to the total basic reproduction number, $\mathcal{R}_{0}$. $\square$
5.3 Sensitivity Analysis and Numerical Results
Remark 5.1 (Parameter Influence Hierarchy)
Analytical and PRCC sensitivity analyses establish the following parameter influence hierarchy on $\mathcal{R}_{0}$:
$\text{High Positive: }\beta>\sigma>\eta_{d}$
$\text{Moderate Positive: }p>\eta_{a}$
$\text{High Negative: }\varepsilon$
$\text{Moderate Negative: }\gamma_{s}>v$
$\text{Low Negative: }\mu_{d}>\alpha>h_{s}$
Classification follows PRCC thresholds: highly influential ($|\text{PRCC}|>0.7$), moderately influential ($0.5<|\text{PRCC}|\leq 0.7$), and less influential ($|\text{PRCC}|\leq 0.5$).
Incorporating targeted integrated strategies aimed at reducing transmission ($\beta$), targeted rapid detection ($\sigma$), and safe burial ($\eta_{d}$) achieve an 85% reduction in outbreaks according to numerical simulations which is a significant increase from vaccination-only strategies that only attain a 45% reduction.
(a) Threshold dynamics verification
(b) Parameter sensitivity analysis
(c) PRCC ranking results
Figure 5: Numerical simulations validating theoretical findings and quantifying parameter impacts
Table 4: Evidence-based intervention prioritization from sensitivity analysis
Priority Interventions Impact
Primary Contact reduction, Safe burial, Early detection 60-85% $\mathcal{R}_{0}$ reduction
Secondary High-efficacy vaccination, Treatment access 25-45% $\mathcal{R}_{0}$ reduction
Tertiary Mass vaccination, Hospital capacity 10-20% $\mathcal{R}_{0}$ reduction
Figure 6: Contours of parameter sensitivity: (a) $\beta,\sigma$ (most sensitive), (b) $\eta_{d},p$ (intermediate), (c) $\varepsilon,\gamma_{s}$ (least sensitive), (d) $\eta_{a},v$ (mixed sensitivity)
The results further affirm the developed theory by demonstrating very strong threshold behavior at $\mathcal{R}_{0}=1$. Such confirmation also supports the parameter sweep results which assert the dominance of the transmission parameters $\beta$, $\sigma$, and $\eta_{d}$, in the dynamics of the outbreaks. Contour plots illustrate the nonlinear interactions which add to the ever-growing complexity of the transmission dynamics of the Ebola virus.
The primary community-based switch strategy employed to control the spread of the virus focuses on the interruption of transmission. Thus, the safe burial practices endorsed by the World Health Organization are critical in a community’s efforts to disturb transmission. Evidence of memory in the Ebola virus, captured by the sophisticated model elements, is indicated by the strong effect of the fractional order $\alpha$. Such results, in turn, offer the mathematical foundation for controlled community-centric Ebola response which is often lacking. This model control will inform the aggressive response necessary for targeted intervention strategies to control the Ebola virus.
@@ 6 Optimal Control Analysis
6.1 Problem Formulation
In this section, we delineate four time-dependent control variables for a complete management of the Ebola virus. Control variables: $u_{1}$ protection [0,0.8], $u_{2}$ vaccination [0,0.15], $u_{3}$ treatment [0,0.8], $u_{4}$ burial [0,0.5] (day-1). The controlled fractional-order system now becomes:
@@ 6.1 Problem Formulation
${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R.$
with modified force of infection $\lambda(t)=\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)/N$. The objective functional minimizes total cost over $[0,T]$:
$J(\mathbf{u})=\int_{0}^{T}\left[\sum_{i=1}^{4}A_{i}X_{i}(t)+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}(t)\right]dt$ (25)
Where $X=[I_{s},I_{a},H,D]$ indicates the infectious compartments, diseased burden costs are denoted by $A_{i}>0$, and the costs related to the implementation of controls are denoted as $B_{j}>0$ in quadratic form.
We look to find the optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that fulfill the requirements of:
$J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u})$ (26)
for the controlled system (24), where the admissible set is:
$\mathcal{U}=\left\{\mathbf{u}\in L^{\infty}(0,T):0\leq u_{i}(t)\leq u_{i}^{\max}\right\}$ (27)
6.2 Necessary Conditions for Optimality
Theorem 6.1 (Existence of Optimal Controls)
Let $J(\mathbf{u})$ be the objective functional relative to the controlled fractional-order system (24) with controllable initial conditions $\mathbf{X}(0)\in\mathbb{R}^{8}_{\geq 0}$, and with $\mathbf{u}\in\mathcal{U}$ being the admissible controls. An optimal control pair $(\mathbf{X}^{*},\mathbf{u}^{*})$ such that $J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u})$ exists.
Proof.  Existence has been demonstrated by verifying the fundamental conditions in [56]. The state space $\mathbb{R}^{8}_{\geq 0}$ is closed and the control set $\mathcal{U}$ is closed, convex and bounded (as per its definition). The system dynamics $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ has the required properties as follows:
The system dynamics $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ are continuous in $(t,\mathbf{X},\mathbf{u})$, is bounded such that $\|\mathbf{f}\|\leq C(1+\|\mathbf{X}\|)$, and is Lipschitz continuous $\|\mathbf{f}(\mathbf{X}_{1})-\mathbf{f}(\mathbf{X}_{2})\|\leq L\|\mathbf{X}_{1}-\mathbf{X}_{2}\|.$ The cost integrand $\mathcal{L}(t,\mathbf{X},\mathbf{u})$ exhibits convexity in $\mathbf{u}$ through quadratic terms and coercivity $\mathcal{L}\geq\frac{1}{2}\min B_{j}\|\mathbf{u}\|^{2}-K$.
The inequality $^{C}D_{t}^{\alpha}N(t)\leq\Lambda-\mu N$ leads to population boundedness, ensuring that $N(t)\leq\max\{N(0),\Lambda/\mu\}$, and thus the state trajectories are bounded. Together, these conditions guarantee the existence of an optimal control pair. $\square$
6.3 Pontryagin’s Maximum Principle for Fractional Systems
@@
Assuming that $(\mathbf{X}^{*},\mathbf{u}^{*})$ is optimal for the control problem, there exist absolutely continuous adjoint variables satisfying the system of fractional differential equations:
@@
${}^{C}D_{t}^{\alpha}\lambda_{S}$ $=-\frac{\partial H}{\partial S},$ ${}^{C}D_{t}^{\alpha}\lambda_{V}$ $=-\frac{\partial H}{\partial V},$ ${}^{C}D_{t}^{\alpha}\lambda_{E}$ $=-\frac{\partial H}{\partial E},$
${}^{C}D_{t}^{\alpha}\lambda_{I_{s}}$ $=-\frac{\partial H}{\partial I_{s}},$ ${}^{C}D_{t}^{\alpha}\lambda_{I_{a}}$ $=-\frac{\partial H}{\partial I_{a}},$ ${}^{C}D_{t}^{\alpha}\lambda_{H}$ $=-\frac{\partial H}{\partial H},$
${}^{C}D_{t}^{\alpha}\lambda_{D}$ $=-\frac{\partial H}{\partial D},$ ${}^{C}D_{t}^{\alpha}\lambda_{R}$ $=-\frac{\partial H}{\partial R},$
The optimal control $\mathbf{u}^{*}(t)$ is described by
@@
supplemented with the transversality condition
$\bm{\lambda}(T)=\mathbf{0},$
with $\bm{\lambda}=(\lambda_{S},\lambda_{V},\lambda_{E},\lambda_{I_{s}},\lambda_{I_{a}},\lambda_{H},\lambda_{D},\lambda_{R})^{\top}$.
Proof.  The proof follows the variational approach for fractional optimal control systems. Consider a perturbation $\mathbf{u}^{\epsilon}=\mathbf{u}^{*}+\epsilon\mathbf{v}$ with $\mathbf{v}\in L^{\infty}([0,T],\mathbb{R}^{4})$ such that $\mathbf{u}^{\epsilon}\in\mathcal{U}$ for sufficiently small $\epsilon>0$. Let $\mathbf{X}^{\epsilon}$ denote the corresponding state trajectory.
The first variation of the objective functional is given by:
where $\delta\mathbf{X}$ satisfies the linearized state equation derived from the controlled system. We have defined adjoint variables $\bm{\lambda}$ which are defined by the following fractional differential equations:
$^{C}D_{t}^{\alpha}\bm{\lambda}=-\frac{\partial H}{\partial\mathbf{X}},\quad\bm{\lambda}(T)=\mathbf{0}.$
Using integration by parts for Caputo fractional derivatives leads to:
$\delta J=\int_{0}^{T}\frac{\partial H}{\partial\mathbf{u}}\cdot\mathbf{v}dt.$
For the problem to be optimally solved, it is necessary to have $\delta J\geq 0$ for every admissible variations $\mathbf{v}$, which leads to:
$\frac{\partial H}{\partial\mathbf{u}}\cdot\mathbf{v}\geq 0\quad\text{a.e. on }[0,T].$
From this inequality, we derive the pointwise minimization condition. Consolidating the minimization condition, we have:
$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}H(t,\mathbf{X}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{a.e.}$
We have thus completed the derivation of necessary optimality conditions. $\square$
6.4 Optimal Control Characterization
@@
The optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that minimize the Hamiltonian are characterized by:
$u_{1}^{*}(t)$ $=\min\left\{u_{1}^{\max},\max\left\{0,\frac{[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)]\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{B_{1}N}\right\}\right\},$
$u_{2}^{*}(t)$ $=\min\left\{u_{2}^{\max},\max\left\{0,\frac{(\lambda_{S}-\lambda_{V})S}{B_{2}}\right\}\right\},$
$u_{3}^{*}(t)$ $=\min\left\{u_{3}^{\max},\max\left\{0,\frac{(\lambda_{I_{s}}-\lambda_{H})I_{s}}{B_{3}}\right\}\right\},$
$u_{4}^{*}(t)$ $=\min\left\{u_{4}^{\max},\max\left\{0,\frac{\lambda_{D}D}{B_{4}}\right\}\right\},$
where all state and adjoint variables follow the optimal trajectory.
Proof.  The characterization derives from solving the first-order optimality conditions $\partial H/\partial u_{i}=0$ for each control variable. For personal protection control $u_{1}^{*}$, the optimality condition yields:
$\frac{\partial H}{\partial u_{1}}=B_{1}u_{1}+\left[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-(S+(1-\varepsilon)V)\lambda_{E}\right]\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}=0.$
Solving algebraically provides the unconstrained optimum:
@@
The last description arises from the projection onto the last interval $0,u_{1}^{\max}]$. For vaccination control $u_{2}^{*}$, the condition:
$\frac{\partial H}{\partial u_{2}}=B_{2}u_{2}+(\lambda_{V}-\lambda_{S})S=0$
gives $u_{2}=(\lambda_{S}-\lambda_{V})S/B_{2}$, with projection ensuring admissibility. For treatment control $u_{3}^{*}$, the optimality condition:
$\frac{\partial H}{\partial u_{3}}=B_{3}u_{3}+(\lambda_{H}-\lambda_{I_{s}})I_{s}=0$
yields $u_{3}=(\lambda_{I_{s}}-\lambda_{H})I_{s}/B_{3}$. For burial control $u_{4}^{*}$, the condition:
$\frac{\partial H}{\partial u_{4}}=B_{4}u_{4}-\lambda_{D}D=0$
provides $u_{4}=\lambda_{D}D/B_{4}$. The satisfying all controls within set projection established within the admis round case the set $\mathcal{U}$ which fulfills the distinctive classification. $\square$
6.5 Adjoint System Specification

### v3

@@ 4.9 GAS of Endemic Equilibrium
The fractional-order Ebola model (2) undergoes a forward bifurcation at $\mathcal{R}_{0}=1$. This bifurcation is characterized by the absence of backward bifurcation and the nonexistence of multiple endemic equilibria.
4.10 Sensitivity Analysis Results
Figure 3: Sensitivity analysis of parameters on basic reproduction number $\mathcal{R}_{0}$ for the Ebola model. Red bars indicate negative sensitivity (parameters that decrease $\mathcal{R}_{0}$), blue bars indicate positive sensitivity (parameters that increase $\mathcal{R}_{0}$).
Table 2: Sensitivity analysis of basic reproduction number ($\mathcal{R}_{0}$) for the Ebola model
Parameter $\Upsilon$ Effect Parameter $\Upsilon$ Effect
$\beta$ +1.0000 High + $\varepsilon$ -0.4123 Medium –
$\sigma$ +0.6250 High + $p$ +0.3752 Medium +
$\eta_{d}$ +0.4678 High + $v$ -0.3561 Medium –
$\eta_{a}$ +0.3117 Medium + $h_{s}$ -0.1987 Low –
$\gamma_{s}$ -0.2541 Medium – $\gamma_{a}$ -0.1876 Low –
$\delta_{s}$ +0.2345 Medium + $\alpha$ -0.1250 Low –
$\omega$ +0.0942 Low + $\mu$ -0.0214 Very low –
$\delta_{h}$ +0.0783 Very low + $\mu_{d}$ -0.0187 Very low –
$\Lambda$ +0.0000 No effect
The sensitivity analysis results, presented in Table 2, reveal the normalized sensitivity indices of model parameters with respect to $\mathcal{R}_{0}$. A positive index indicates that $\mathcal{R}_{0}$ increases with the parameter, while a negative index corresponds to a decrease in $\mathcal{R}_{0}$. Parameters with negative sensitivity indices ($\varepsilon$, $v$, $\gamma_{s}$, $h_{s}$, $\gamma_{a}$, $\alpha$, $\mu$, $\mu_{d}$) reduce $\mathcal{R}_{0}$ when increased, thereby suppressing disease transmission. Conversely, parameters with positive indices ($\beta$, $\sigma$, $\eta_{d}$, $p$, $\eta_{a}$, $\delta_{s}$, $\omega$, $\delta_{h}$) enhance $\mathcal{R}_{0}$ when increased, leading to increased transmission potential. The transmission rate $\beta$ exhibits the highest positive sensitivity index (+1.0000), indicating that $\mathcal{R}_{0}$ is directly proportional to $\beta$. The significance of contact reduction measures like social distancing and personal protection gear is highlighted by this. The crucial importance of secure burial practices provides for controlling outbreaks is demonstrated by the dead transmission rate $\eta_{d}$ (index +0.4678). Even at moderate coverage levels, vaccinations may reduce transmission, as indicated by the negative sensitivity (-0.4123) of vaccine efficacy $\varepsilon$. Interestingly, the hospitalization rate \$h_s\$ has a negative sensitivity index (-0.1987), indicating that through isolating infectious persons, expanding hospital capacity lowers transmission. The fractional order $\alpha$ displays low sensitivity (-0.1250), suggesting that decreased transmission potential in the fractional-order model is an outcome of memory effects. The recruitment rate $\Lambda$ has zero sensitivity, as expected from the structure of $\mathcal{R}_{0}$, indicating that population turnover does not directly affect the reproduction number in this model formulation. Figure 3 displays a tornado plot of the sensitivity indices, ranking parameters by their influence on $\mathcal{R}_{0}$. The dominance of $\beta$, $\sigma$, and $\eta_{d}$ as the top three positive-sensitivity parameters reinforces the importance of transmission reduction, rapid case detection, and safe burial practices in Ebola control strategies.
@@ 5 Optimal Control Analysis
In this section, we introduce four time-dependent control variables to manage Ebola transmission effectively. The controls are personal protection ($u_{1}$, range [0,0.8]), vaccination ($u_{2}$, range [0,0.15]), treatment ($u_{3}$, range [0,0.8]), and safe burial ($u_{4}$, range [0,0.5]) (all in units day-1).
The controlled fractional-order system is given by
@@
${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R,$
with the modified force of infection $\lambda(t)=\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)/N$. The objective functional to be minimized over the time horizon $[0,T]$ is
$J(\mathbf{u})=\int_{0}^{T}\left[\sum_{i=1}^{4}A_{i}X_{i}(t)+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}(t)\right]dt,$ (49)
where $X=[I_{s},I_{a},H,D]$ represents the infectious compartments, $A_{i}>0$ are the cost weights associated with the disease burden, and $B_{j}>0$ are the quadratic cost coefficients for implementing the controls. We seek the optimal controls $\mathbf{u}^{*}=(u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*})$ that satisfy
$J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u}),$ (50)
subject to the dynamics (48), where the admissible control set is
$\mathcal{U}=\left\{\mathbf{u}\in L^{\infty}(0,T):0\leq u_{i}(t)\leq u_{i}^{\max}\right\}.$ (51)
@@
Let the fractional-order controlled system (48) with initial conditions $\mathbf{X}(0)\in\mathbb{R}^{8}_{\geq 0}$ and admissible control set $\mathcal{U}$ satisfy
  1. (i)
    The control set $\mathcal{U}$ is closed, convex, and bounded in $L^{\infty}(0,T)$.
  2. (ii)
    The right-hand side $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ of (48) is continuous in $(t,\mathbf{X},\mathbf{u})$, measurable in $t$, and satisfies as
    • Linear growth $\|\mathbf{f}(t,\mathbf{X},\mathbf{u})\|\leq c_{1}+c_{2}\|\mathbf{X}\|+c_{3}\|\mathbf{u}\|$ for constants $c_{1},c_{2},c_{3}>0$.
    • Lipschitz continuity $\|\mathbf{f}(t,\mathbf{X}_{1},\mathbf{u})-\mathbf{f}(t,\mathbf{X}_{2},\mathbf{u})\|\leq L\|\mathbf{X}_{1}-\mathbf{X}_{2}\|$ for some $L>0$.
  3. (iii)
    The integrand $\mathcal{L}(t,\mathbf{X},\mathbf{u})=\sum_{i=1}^{4}A_{i}X_{i}(t)+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}(t)$ of the objective functional (49) is convex in $\mathbf{u}$ and satisfies the coercivity condition $\mathcal{L}(t,\mathbf{X},\mathbf{u})\geq c_{4}\|\mathbf{u}\|^{2}-c_{5}$ for constants $c_{4}>0$, $c_{5}\geq 0$.
Then there exists an optimal control pair $(\mathbf{X}^{*},\mathbf{u}^{*})$ that minimizes $J(\mathbf{u})$ subject to (48).
Proof.  We verify the conditions of the Filippov-Cesari existence theorem for fractional optimal control problems [55, 56]. By definition, $\mathcal{U}=\{\mathbf{u}\in L^{\infty}(0,T):0\leq u_{i}(t)\leq u_{i}^{\max}\}$ is closed, convex, and bounded in $L^{\infty}(0,T)$.
The right-hand side $\mathbf{f}$ of (48) is polynomial in $\mathbf{X}$ and $\mathbf{u}$, hence continuous and measurable. For the linear growth condition, note that all terms are at most quadratic (products like $(1-u_{1})\lambda S$ where $\lambda$ is linear in $I_{s},I_{a},D$). Since $N(t)\geq\mu S^{*}>0$ (by Theorem 1) and the population is bounded by $\max\{N(0),\Lambda/\mu\}$, there exists $M>0$ such that $\|\mathbf{f}(t,\mathbf{X},\mathbf{u})\|\leq M(1+\|\mathbf{X}\|+\|\mathbf{u}\|)$. Lipschitz continuity follows from the boundedness of $\mathbf{u}$ and the fact that all partial derivatives $\partial\mathbf{f}/\partial\mathbf{X}$ are bounded on bounded sets.
The integrand $\mathcal{L}$ is quadratic in $\mathbf{u}$, hence convex in $\mathbf{u}$. Moreover, with $B_{\min}=\min\{B_{1},B_{2},B_{3},B_{4}\}>0$, we have $\mathcal{L}(t,\mathbf{X},\mathbf{u})\geq\frac{1}{2}B_{\min}\|\mathbf{u}\|^{2}$, satisfying the coercivity condition.
From the population dynamics $^{C}D_{t}^{\alpha}N\leq\Lambda-\mu N$, we have $N(t)\leq\max\{N(0),\Lambda/\mu\}$. Since all compartments are nonnegative and $N=S+V+E+I_{s}+I_{a}+H+D+R$, each compartment is bounded by this maximum.
All conditions of Theorem 2.1 in [55] (existence theorem for fractional optimal control problems) are satisfied. Therefore, there exists an optimal control pair $(\mathbf{X}^{*},\mathbf{u}^{*})$ minimizing $J(\mathbf{u})$. $\square$
Theorem 5.2
@@
Let $(\mathbf{X}^{*},\mathbf{u}^{*})$ be an optimal solution for the control problem defined by the objective functional (49) subject to the fractional-order system (48). Then there exist absolutely continuous adjoint variables $\bm{\lambda}=(\lambda_{S},\lambda_{V},\lambda_{E},\lambda_{I_{s}},\lambda_{I_{a}},\lambda_{H},\lambda_{D},\lambda_{R})^{\top}$ satisfying the fractional adjoint equations
${}^{C}D_{t}^{\alpha}\lambda_{S}$ $=-\frac{\partial\mathcal{H}}{\partial S},\quad{}^{C}D_{t}^{\alpha}\lambda_{V}=-\frac{\partial\mathcal{H}}{\partial V},\quad{}^{C}D_{t}^{\alpha}\lambda_{E}=-\frac{\partial\mathcal{H}}{\partial E},$ (52)
${}^{C}D_{t}^{\alpha}\lambda_{I_{s}}$ $=-\frac{\partial\mathcal{H}}{\partial I_{s}},\quad{}^{C}D_{t}^{\alpha}\lambda_{I_{a}}=-\frac{\partial\mathcal{H}}{\partial I_{a}},\quad{}^{C}D_{t}^{\alpha}\lambda_{H}=-\frac{\partial\mathcal{H}}{\partial H},$ (53)
${}^{C}D_{t}^{\alpha}\lambda_{D}$ $=-\frac{\partial\mathcal{H}}{\partial D},\quad{}^{C}D_{t}^{\alpha}\lambda_{R}=-\frac{\partial\mathcal{H}}{\partial R},$ (54)
where the Hamiltonian $\mathcal{H}$ is defined as
$\mathcal{H}(t,\mathbf{X},\mathbf{u},\bm{\lambda})=\mathcal{L}(t,\mathbf{X},\mathbf{u})+\bm{\lambda}^{\top}\mathbf{f}(t,\mathbf{X},\mathbf{u}),$
with $\mathcal{L}(t,\mathbf{X},\mathbf{u})=\sum_{i=1}^{4}A_{i}X_{i}+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}$ and $\mathbf{f}$ denoting the right-hand side of (48).
The optimal control $\mathbf{u}^{*}(t)$ satisfies the minimization condition
$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}\mathcal{H}(t,\mathbf{X}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{almost everywhere on }[0,T],$ (55)
with the transversality condition
$\bm{\lambda}(T)=\mathbf{0}.$ (56)
Proof.  The proof follows the variational approach for fractional optimal control systems. Let $(\mathbf{X}^{*},\mathbf{u}^{*})$ be an optimal pair. Consider a perturbation $\mathbf{u}^{\epsilon}=\mathbf{u}^{*}+\epsilon\mathbf{v}$, where $\mathbf{v}\in L^{\infty}([0,T],\mathbb{R}^{4})$ satisfies $\mathbf{u}^{\epsilon}\in\mathcal{U}$ for sufficiently small $\epsilon>0$. Let $\mathbf{X}^{\epsilon}$ denote the corresponding state trajectory.
The first variation of the objective functional is given by
where $\delta\mathbf{X}$ satisfies the linearized state equation derived from (48).
Define the adjoint variables $\bm{\lambda}$ through the fractional boundary value problem
$^{C}D_{t}^{\alpha}\bm{\lambda}=-\frac{\partial\mathcal{H}}{\partial\mathbf{X}},\quad\bm{\lambda}(T)=\mathbf{0}.$ (57)
Applying the fractional integration by parts formula for Caputo derivatives to the term involving $\delta\mathbf{X}$ yields
@@
$\int_{0}^{T}\frac{\partial\mathcal{L}}{\partial\mathbf{X}}\cdot\delta\mathbf{X}\,dt=\int_{0}^{T}\bm{\lambda}\cdot\left[\frac{d}{dt}\delta\mathbf{X}\right]\,dt=-\int_{0}^{T}\left[{}^{C}D_{t}^{\alpha}\bm{\lambda}\right]\cdot\delta\mathbf{X}\,dt.$
Substituting the adjoint equation (57) and simplifying gives
$\delta J=\int_{0}^{T}\left[\frac{\partial\mathcal{H}}{\partial\mathbf{u}}\right]\cdot\mathbf{v}\,dt.$
For optimality, we require $\delta J\geq 0$ for all admissible variations $\mathbf{v}$, which implies the pointwise condition
$\frac{\partial\mathcal{H}}{\partial\mathbf{u}}\cdot\mathbf{v}\geq 0\quad\text{almost everywhere on }[0,T].$
This inequality leads to the minimization condition
@@
completing the derivation of the necessary optimality conditions. $\square$
@@
The optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that minimize the Hamiltonian $\mathcal{H}$ are given by
$u_{1}^{*}(t)$ $=\min\left\{u_{1}^{\max},\max\left\{0,\frac{[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)]\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{B_{1}N}\right\}\right\},$ (58)
$u_{2}^{*}(t)$ $=\min\left\{u_{2}^{\max},\max\left\{0,\frac{(\lambda_{S}-\lambda_{V})S}{B_{2}}\right\}\right\},$ (59)
$u_{3}^{*}(t)$ $=\min\left\{u_{3}^{\max},\max\left\{0,\frac{(\lambda_{I_{s}}-\lambda_{H})I_{s}}{B_{3}}\right\}\right\},$ (60)
$u_{4}^{*}(t)$ $=\min\left\{u_{4}^{\max},\max\left\{0,\frac{\lambda_{D}D}{B_{4}}\right\}\right\},$ (61)
where all state and adjoint variables follow their optimal trajectories.
Proof.  The optimal controls are derived by solving the first-order optimality conditions $\partial\mathcal{H}/\partial u_{i}=0$ for each control variable, then projecting the solutions onto the admissible interval $0,u_{i}^{\max}]$.
For personal protection control $u_{1}$, the optimality condition provides
$\frac{\partial\mathcal{H}}{\partial u_{1}}=B_{1}u_{1}+\left[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)\right]\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}=0.$
Solving for $u_{1}$ gives the unconstrained optimum
@@
Projecting this expression onto the admissible interval $0,u_{1}^{\max}]$ yields the optimal control (58).
For vaccination control $u_{2}$, the optimality condition is
$\frac{\partial\mathcal{H}}{\partial u_{2}}=B_{2}u_{2}+(\lambda_{V}-\lambda_{S})S=0.$
Solving gives $u_{2}=(\lambda_{S}-\lambda_{V})S/B_{2}$, and after projection we obtain (59).
For treatment control $u_{3}$, we have
$\frac{\partial\mathcal{H}}{\partial u_{3}}=B_{3}u_{3}+(\lambda_{H}-\lambda_{I_{s}})I_{s}=0,$
which yields $u_{3}=(\lambda_{I_{s}}-\lambda_{H})I_{s}/B_{3}$. Projection gives (60).
For safe burial control $u_{4}$, the condition
$\frac{\partial\mathcal{H}}{\partial u_{4}}=B_{4}u_{4}-\lambda_{D}D=0$
gives $u_{4}=\lambda_{D}D/B_{4}$. Projection onto $0,u_{4}^{\max}]$ yields (61).
Each optimal control is thus characterized by projecting the solution of the first-order optimality condition onto the admissible control set $\mathcal{U}$, ensuring that all constraints $0\leq u_{i}(t)\leq u_{i}^{\max}$ are satisfied. $\square$